### v1

# supplement

Abstract
We investigate two transformations within the framework of compositional data analysis for forecasting the age distribution of death counts. Drawing on age-specific period life-table death counts from 24 countries in the Human Mortality Database, we assess and compare the point and interval forecast accuracy of the two transformations. Enhancing the forecast accuracy of period life-table death counts holds significant value for demographers, who rely on such forecasts to estimate survival probabilities and life expectancy, and for actuaries, who use them to price temporary annuities across various entry ages and maturities. While our primary focus is on temporary annuities, we also consider long-term contracts that, particularly at higher entry ages, approximate lifetime annuities.
Keywords: centered log-ratio transformation; cumulative distribution function transformation; life-table death counts; principal component analysis; single-premium temporary annuity
1 Introduction
Actuaries and demographers have long been engaged in mortality modeling and forecasting, dating back to the early 20th century, as a response to the financial challenges posed by rapid mortality improvements. These improvements affected the accuracy of annuity calculations and the sustainability of government pension systems (Pollard, 1987). Forecasting life tables for annuitants was a key topic at the 5th International Congress of Actuaries in Berlin in 1906 (Cramér & Wold, 1935). In the field of statistical demography, numerous researchers have proposed innovative approaches to modeling and forecasting mortality by age, sex, and cause, across time and space. Special emphasis has been placed on statistical time-series extrapolation models (see, e.g., Booth, 2006; Booth & Tickle, 2008; Cairns et al., 2008; Shang et al., 2011; Basellini et al., 2023, for comprehensive reviews). Within mortality modeling, at least three instruments are commonly used: age-specific mortality rates, survival probabilities, and the age distribution of death counts. Bergeron-Boucher et al. (2019) provide a thorough analysis of these mortality instruments, highlighting their respective advantages and limitations. All three instruments can be derived from one another using standard life-table relationships (see Preston et al., 2001, Chapter 3). Computationally, the conversion among these instruments can be achieved using the LifeTable function in the MortalityLaws package (Pascariu, 2025).
Among the three mortality instruments, we focus on life-table death counts and model their evolution over time. These counts reveal a clear shifting pattern, with deaths progressively moving from younger to older ages. Between cohort and period life-table death counts, we choose to work with the latter, as period life tables are not only complete but also reflect the prevailing mortality conditions within a given time frame (see also Oeppen, 2008; Bergeron-Boucher et al., 2017). A time series of period life-table death counts enables the analysis of temporal changes in key longevity indicators, such as the modal age at death (see, e.g., Canudas-Romo, 2010), as well as measures of lifespan variability, including the standard deviation, interquartile range, and Gini coefficient (see, e.g., Vaupel et al., 2011; van Raalte & Caswell, 2013).
In demography, Oeppen (2008) introduced the use of principal component analysis to model and forecast the age distribution of deaths within a compositional data analysis framework, treating age-specific life-table death counts ($d_{x}$) as compositional data. As with all compositional data, these counts are subject to inherent constraints, typically bounded between zero and a fixed upper limit. Consequently, the sample space of compositional data is a simplex, defined as
$\mathbb{S}^{D-1}=\left\{(d_{1},\dots,d_{D})^{\top},\;d_{x}\geq 0,\;\sum^{D}_{x=1}d_{x}=c\right\},$
where $\mathbb{S}$ denotes a simplex, which is a $(D-1)$-dimensional subset of real-valued space $R^{D-1}$, denotes vector transpose, and $c$ is a fixed constant, set typically to one (portions) (Scealy et al., 2015), $10^{6}$ parts per million (Scealy et al., 2015), or $10^{5}$ life-table death counts.
Due to the non-negativity and summability constraints inherent in compositional data, conventional linear techniques may be inappropriate, as they do not account for the unique variance-covariance structure of such data. A common strategy is to first transform the data to remove the summability constraint, thereby enabling the application of linear methods. One widely used approach is the family of log-ratio transformations, particularly the centered log-ratio (CLR) transformation (see, e.g., Aitchison, 1986). An alternative is the cumulative distribution function (CDF) transformation, which leverages monotonicity to effectively handle the presence of zero values (see Shang & Haberman, 2025).
This paper makes two key contributions. First, we evaluate and compare the point forecast accuracy of the CLR and CDF transformations, using data from 24 countries in the Human Mortality Database (2025), with time series starting in or before 1950. Second, while the existing literature has largely focused on point forecasts, we introduce a computationally efficient approach for constructing pointwise prediction intervals and compare the interval forecast accuracy between the CLR and CDF transformations.
The remainder of this paper is structured as follows. Section 2 describes the multi-country data sets used in the analysis. Section 3 revisits the CLR and CDF transformations and introduces a forecasting method based on principal component analysis. In Section 4, we assess point forecast accuracy using the Kullback-Leibler and Jensen-Shannon divergences. For evaluating interval forecast accuracy, we consider the empirical coverage probability and its absolute deviation from the nominal coverage level, as well as the interval score proposed by Gneiting & Raftery (2007). Section 5 applies the CDF transformation to estimate the single-premium temporary immediate annuity prices for various entry ages and maturities for female and male populations in 24 countries. Section 6 concludes with a summary and suggestions for future extensions of the proposed methodology.
2 Age distribution of death counts
The data sets used in this study are sourced from the Human Mortality Database (2025). We consider single-year ages ranging from 0 to 110. Out of 41 predominantly developed countries, we select 24 countries for which life-table death counts are available from 1950 or earlier. This results in 48 sex-specific populations used for analysis and comparison. The selected countries and their corresponding sample periods are listed in Table 1.
Table 1: Countries and their respective sample periods. For the annuity price calculations, we use the official interest rate as of 18 May 2025.
Country Code Period Interest rate Country Code Period Interest rate
Australia AUS 1921–2021 4.10% Italy ITA 1872–2022 2.25%
Austria AUT 1947–2023 2.25% Japan JPN 1947–2023 0.50%
Belgium BEL 1919–2023 2.25% Netherlands NLD 1850–2022 2.25%
Bulgaria BGR 1947–2021 2.24% Norway NOR 1846–2023 4.50%
Canada CAN 1921–2022 4.50% New Zealand NZ 1948–2021 5.50%
Czech CZE 1950–2021 3.50% Portugal PRT 1940–2023 2.25%
Denmark DEN 1835–2024 2.25% Slovakia SVK 1950–2019 2.25%
Finland FIN 1878–2023 2.25% Spain SPA 1908–2023 2.25%
France FRA 1816–2022 2.25% Sweden SWE 1751–2023 4.00%
Hungary HUN 1950–2020 6.50% Switzerland CHE 1876–2023 1.75%
Iceland ICE 1838–2023 7.75% United Kingdom UK 1922–2022 5.25%
Ireland IRE 1950–2022 2.25% United States USA 1933–2023 5.25%
The life-table death counts are non-negative and sum to a radix (i.e., a population experiencing 100,000 births annually) for each year. For the life-table death counts, there are 111 ages, and these are ages 0, 1, $\dots$, 109, 110+. Due to rounding, there may be zero counts for higher ages at some years. To rectify this problem, we work with the probability of dying (i.e., $q_{x}$) and the life-table radix to recalculate our estimated death counts (up to six decimal places). In doing so, we obtain more detailed death counts than the ones reported.
To understand the main features of the data, Figure 1 displays rainbow plots of the female and male age-specific life-table death counts in Australia in a single-year group from 1921 to 2021. The time ordering of the curves follows the colour order of a rainbow, where data from the distant past are shown in red while the more recent data are shown in purple (see Hyndman & Shang, 2010, for other examples). Both figures demonstrate a decreasing trend in infant death counts and a typical negatively skewed distribution, where the peaks gradually shift to higher ages for both sexes. This shift is a source of longevity risk, which creates a major issue for insurers and government pension funds, especially in the selling and risk management of annuity products (see Denuit et al., 2007, for a discussion). Moreover, the spread of the distribution indicates lifespan variability. A decrease in variability over time can be observed directly and can be measured, with interquartile range or Gini coefficient (see, e.g., Wilmoth & Horiuchi, 1999; van Raalte & Caswell, 2013; Debón et al., 2017). With the decrease in infant deaths and increase in mortality at older ages, the spread of the distribution has been reduced.
Figure 1: Rainbow plots of age-specific life-table death counts from 1921 to 2021 in a single-year group. The oldest years are shown in red, with the most recent years in violet. Curves are ordered chronologically according to the colours of the rainbow.
3 Transformation
Let the age-specific life-table death counts be denoted by $d^{s}_{t,i}$ or $d^{s}_{t,x}$, where $t$ denotes a year, $s$ represents females or males, and $i$ or $x$ denotes an age. For each year $t$, the life-table death counts sum to a radix $10^{5}$.
3.1 Centered log-ratio transformation
The CLR transformation presents a mapping from the simplex to a hyperplane in Euclidean space, where principal component analysis can be applied. The transformation consists of the following steps:
  1. 1)
    Compute the geometric mean function, estimated by a simple average
    $\alpha^{s}_{n,x}=\exp\Big{(}\frac{1}{n}\sum^{n}_{t=1}\ln d^{s}_{t,x}\Big{)}.$
    We compute the centered data by dividing each death count by its geometric mean
    $\zeta^{s}_{t,x}=\frac{d^{s}_{t,x}}{\alpha_{n,x}^{s}}.$
  2. 2)
    By taking the natural logarithm, we obtain
    $\beta_{t,x}^{s}$ $=\ln\zeta^{s}_{t,x}$
    $=\ln d_{t,x}^{s}-\ln\alpha_{n,x}^{s}$
    $=\ln d_{t,x}^{s}-\frac{1}{n}\sum^{n}_{t=1}\ln d_{t,x}^{s},$
    where $\bm{\beta}_{x}^{s}=(\beta_{1,x}^{s},\dots,\beta_{n,x}^{s})^{\top}$ denotes a set of unconstrained functional time series.
  3. 3)
    By applying eigendecomposition to $\bm{\beta}_{x}^{s}$, we can express
    $\beta_{t,x}^{s}=\sum^{L_{s}}_{\ell=1}\gamma_{t,\ell}\phi_{\ell,x}^{s}+\omega_{t,x}^{s},$
    where $\phi_{\ell,x}^{s}$ is the $\ell$th estimated principal component for age $x$, $\gamma^{s}_{t,\ell}=\langle\beta_{t,x}^{s},\phi_{\ell,x}^{s}\rangle$ is the $\ell$th estimated principal component score at time $t$, $\omega_{t,x}^{s}$ denotes the model residuals for age $x$ in year $t$, and $L_{s}$ denotes the number of retained principal components. For selecting the number of components $L_{s}$, we implement an eigenvalue ratio criterion of Li et al. (2020), as well as $L_{s}=6$ recommended in Hyndman et al. (2013).
  4. 4)
    By conditioning on the estimated principal components and observed data, the $h$-step-ahead forecast of $\beta_{n+h,x}^{s}$ can be estimated by
    $\widehat{\beta}^{s}_{n+h|n,x}=\sum^{L_{s}}_{\ell=1}\widehat{\gamma}^{s}_{n+h|n,\ell}\phi_{\ell,x}^{s},$
    where $\widehat{\gamma}_{n+h|n,\ell}$ denotes the $h$-step-ahead forecast of the principal component scores.
  5. 5)
    By taking the inverse CLR transformation, we traverse back to the original simplex
    $\widehat{\zeta}_{n+h|n,x}^{s}=\exp^{\widehat{\beta}^{s}_{n+h|n,x}}.$
  6. 6)
    Then, we add back the geometric mean to obtain the forecasts of the life-table death count $d_{n+h,x}^{s}$:
    $\widehat{d}_{n+h|n,x}^{s}=\widehat{\zeta}_{n+h|n,x}^{s}\times\alpha_{n,x}^{s}.$
3.2 Cumulative distribution function transformation
For a time series of CDFs, it is important to scale the life-table death counts to sum to one. We describe the transformation in the following steps:
  1. 1)
    Compute the empirical cumulative distribution function via the cumulative sum,
    $D^{s}_{t,x}=\sum^{x}_{i=1}d^{s}_{t,i},\quad x=1,\dots,111,\quad t=1,\dots,n,$
    where $D^{s}_{t,111}=1$.
  2. 2)
    Perform the logistic transformation,
    $Z^{s}_{t,x}=\text{logit}(D^{s}_{t,x})=\ln\left(\frac{D^{s}_{t,x}}{1-D^{s}_{t,x}}\right),$
    where $\ln(\cdot)$ denotes the natural logarithm.
  3. 3)
    With a set of unconstrained data $\bm{Z}^{s}_{x}=(Z^{s}_{1,x},\dots,Z^{s}_{n,x})^{\top}$, we implement a univariate functional time-series forecasting method. By computing the sample variance, we express a stochastic process $Z^{s}_{t,x}$ as
    $Z^{s}_{t,x}=\sum^{K_{s}}_{k=1}\eta^{s}_{t,k}\psi^{s}_{k,x}+\varepsilon^{s}_{t,x},$
    where $\psi^{s}_{k,x}$ denotes the $k$th principal component for age $x$ and sex $s$, $\eta^{s}_{t,k}=\langle Z^{s}_{t,x},\psi^{s}_{k,x}\rangle$ is the estimated principal component score at time $t$, $\varepsilon^{s}_{t,x}$ denotes the model residual function for sex $s$ in year $t$, and $K_{s}$ denotes the number of retained principal components. We implement an eigenvalue ratio criterion of Li et al. (2020), as well as $K_{s}=6$ recommended in Hyndman et al. (2013).
  4. 4)
    By conditioning on the estimated functional principal components $\bm{\Psi}^{s}_{x}=(\psi_{1,x},\dots,\psi_{K_{s},x})$ and observed data $\bm{Z}_{x}^{s}=(Z_{1,x}^{s},\dots,Z_{n,x}^{s})$, the $h$-step-ahead forecast of $Z_{n+h,x}^{s}$ can be obtained
    $\widehat{Z}_{n+h|n,x}^{s}=\text{E}[Z_{n+h,x}^{s}|\bm{\Psi}^{s}_{x},\bm{Z}_{x}^{s}]=\sum^{K_{s}}_{k=1}\widehat{\eta}_{n+h|n,k}^{s}\psi_{k,x}^{s},$
    where $\widehat{\eta}_{n+h|n,k}^{s}$ denotes the $h$-step-ahead univariate time-series forecast of the principal component scores. Among the univariate time-series methods, we consider the exponential smoothing method of Hyndman et al. (2008) to model and forecast each set of scores.
  5. 5)
    By taking the inverse logit transformation, we obtain
    $\widehat{D}_{n+h|n,x}^{s}=\frac{\exp^{\widehat{Z}_{n+h|n,x}^{s}}}{1+\exp^{\widehat{Z}_{n+h|n,x}^{s}}}.$
  6. 6)
    By taking the first-order differencing, we obtain
    $\widehat{d}^{s}_{n+h|n,x}=\Delta^{x}_{i=1}\widehat{D}^{s}_{n+h|n,i},$
    where $\Delta$ represents the first-order differencing.
4 Forecast accuracy comparison
4.1 Expanding window
We implement an expanding-window scheme, which is commonly used to evaluate model and parameter stability, as well as prediction accuracy over time. The expanding window analysis determines the stability of a model’s parameter estimates and their forecasts over an expanding window through the sample. Since each data set has a different sample period, we keep the last 20 years of data for evaluation and the remaining data for training. For the 20 years of data in the testing sample, we consider one-step-ahead to 20-step-ahead forecasts to assess the model’s performance at the short to medium horizons. Through the expanding-window scheme, we produce 20 one-step-ahead forecasts, 19 two-step-ahead forecasts, $\dots$, and one 20-step-ahead forecast. We compare these forecasts with the holdout data in the testing sample to determine the accuracy of the out-of-sample forecast. In Figure 2, we present a concept diagram.
[…]
Figure 2: A diagram of the expanding-window forecast scheme.
4.2 Point forecast evaluation metrics
Since the age distribution of death counts can be considered a probability density function, we consider two density evaluation metrics. The metrics are discrete Kullback-Leibler divergence (KLD) (Kullback & Leibler, 1951) and Jensen-Shannon divergence (JSD) (Shannon, 1948).
The KLD measures information loss by approximating an unknown density with its approximation. For two probability density functions, denoted by $d_{n+\xi}(u)$ and $\widehat{d}_{n+\xi|n}(u)$, the discrete KLD is defined as
$\text{KLD}(h)=$ $D_{\text{KL}}(d_{n+\xi,x}||\widehat{d}_{n+\xi|n,x})+D_{\text{KL}}(\widehat{d}_{n+\xi|n,x}||d_{n+\xi,x})$
=\ = $\frac{1}{111\times(21-h)}\sum^{20}_{\xi=h}\sum^{111}_{x=1}d_{n+\xi,x}\cdot(\ln d_{n+\xi,x}-\ln\widehat{d}_{n+\xi|n,x})+$
$\frac{1}{111\times(21-h)}\sum^{20}_{\xi=h}\sum^{111}_{x=1}\widehat{d}_{n+\xi|n,x}\cdot(\ln\widehat{d}_{n+\xi|n,x}-\ln d_{n+\xi,x}),$
where $\xi$ denotes the forecasting period. A feature of the discrete KLD is its non-negativity.
Even though we have a sample of observed densities in the forecasting period, they may not be the actual densities. An alternative is given by the JSD, which can be seen as a symmetric and smoothed version of the KLD. The JSD is defined by
$\text{JSD}(h)=\frac{1}{2}D_{\text{KL}}(d_{n+\xi,x}||\delta_{n+\xi,x})+\frac{1}{2}D_{\text{KL}}(\widehat{d}_{n+\xi|n,x}||\delta_{n+\xi,x}),$
where $\delta_{n+\xi,x}$ measures a common quantity between $d_{n+\xi,x}$ and $\widehat{d}_{n+\xi,x}$. One example of $\delta_{n+\xi,x}$ can be its geometric mean $\delta_{n+\xi,x}=\sqrt{d_{n+\xi,x}\widehat{d}_{n+\xi,x}}$.
4.3 Comparison of point forecast accuracy
Averaging over the point forecast errors obtained from the 24 countries, Figure 3 presents horizon-specific plots to compare the accuracy of the two transformations, as measured by the KLD and JSD. Between the CLR and CDF transformations, the CDF transformation consistently outperforms the CLR transformation. Such superiority is credited to the monotonicity constraint enjoyed by the CDF transformation. For modeling the unconstrained data in the Euclidean space, it is advantageous to use $K=6$ for forecasting. In contrast, the number of components selected by the EVR tends to be fewer than the optimal. Choosing a larger number of components results in only a minor loss, whereas selecting too few components leads to a more substantial loss.
Figure 3: Comparison of the point forecast accuracy, measured by the KLD and JSD, between the CLR and CDF transformations. For each transformation, we implement a principal component analysis where the number of components is selected by the EVR criterion or fitted at $K=6$.
4.4 Construction of pointwise prediction intervals
Using the data in the training sample, we carry out an expanding window approach to obtain the $h$-step-ahead density forecasts in the validation set for $h=1,\dots,19$. For different forecast horizons, we have different numbers of curves in the validation set. For $h=1$, we have 20 years in evaluate the forecast errors; when $h=19$, we have two years to evaluate the residual functions between the holdout data in the validation set and their forecasts. Based on these residuals, we can compute the functional standard deviation, which requires at least two years of data. The forecast errors can be denoted by $\widehat{\epsilon}_{\nu,x}=d_{\nu,x}-\widehat{d}_{\nu,x}$, where $\widehat{d}_{\nu,x}$ denotes the forecasts obtained from the two transformations, and $\nu=1,\dots,(21-h)$ denotes the number of observations in the validation set for $h=1,\dots,19$.
Let us denote $\Gamma_{x}=\text{sd}(\widehat{\epsilon}_{\nu,x})$, where $\text{sd}(\cdot)$ can be pointwise standard deviation. We denote $\bm{\Gamma}=(\Gamma_{1},\dots,\Gamma_{111})^{\top}$ and $\bm{\widehat{\epsilon}}_{\nu}=(\widehat{\epsilon}_{\nu,1},\dots,\widehat{\epsilon}_{\nu,111})^{\top}$. For a level of significance $\alpha$, customarily $\alpha=0.2$ or 0.05, we aim to find $\theta_{\alpha}$ such that approximately $100(1-\alpha)\%$ of the residuals satisfy
$-\theta_{\alpha}\bm{\Gamma}\leq\bm{\widehat{\epsilon}}_{\nu}\leq\theta_{\alpha}\bm{\Gamma},\qquad v=1,\dots,(21-h).$
By the law of large numbers, one may achieve
$\text{Pr}(-\theta_{\alpha}\Gamma_{x}\leq d_{n+h,x}-\widehat{d}_{n+h|n,x}\leq\theta_{\alpha}\Gamma_{x})\approx\frac{1}{21-h}\sum^{21-h}_{\nu=1}\mathds{1}(-\theta_{\alpha}\bm{\Gamma}\leq\bm{\widehat{\epsilon}}_{\nu}\leq\theta_{\alpha}\bm{\Gamma}),$
where $\mathds{1}(\cdot)$ represents the binary indicator function.
To determine the optimal $\theta_{\alpha}$, the samples in the validation set are used to calibrate a prediction interval so that its empirical coverage probability is close to its nominal coverage probability. With the estimated $\theta_{\alpha}$, the $100(1-\alpha)\%$ prediction interval of $d_{n+h,x}$ can be obtained as
$\big{[}\widehat{d}_{n+h|n,x}-\theta_{\alpha}\Gamma_{x},\quad\widehat{d}_{n+h|n,x}+\theta_{\alpha}\Gamma_{x}\big{]}.$
4.5 Interval forecast evaluation metrics
For each year in the forecasting period, the $h$-step-ahead prediction intervals were computed at the $(1-\alpha)$ nominal coverage probability, where $\alpha$ denotes a level of significance. For a chosen $\alpha$, let us denote $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$, as the lower and upper bounds, respectively. We compute the empirical coverage probability, which can be defined as the number of holdout data points between the lower and upper bounds for a given level of significance. It can be expressed as
$\text{ECP}_{\alpha}(h)=1-\frac{\sum^{20}_{\xi=h}\sum^{111}_{x=1}\left[\mathds{1}(d_{n+\xi,x}<\widehat{d}_{n+\xi|n,x}^{\text{lb}})+\mathds{1}(d_{n+\xi,x}>\widehat{d}_{n+\xi|n,x}^{\text{ub}})\right]}{111\times(21-h)},$
where the denominator is the number of curves in the forecasting period.
From the ECP, we evaluate the coverage probability difference (CPD), which is the absolute difference between empirical and nominal coverage probabilities, and use this measure to evaluate the accuracy of the interval forecast. As an absolute measure, the CPD cannot inform us about the undercoverage or overcoverage. However, it eliminates the possibility of the canceling effect, where undercoverage and overcoverage may concurrently occur. The CPD can be expressed as
$\text{CPD}_{\alpha}(h)=|\text{ECP}_{\alpha}(h)-(1-\alpha)|.$
The smaller the value of CPD is, the more accurate the prediction interval one method provides.
The ECP and CPD measure the accuracy of prediction intervals, and neither considers the sharpness of the prediction intervals, i.e., the distance between lower and upper bounds. We also consider the interval score of Gneiting & Raftery (2007), defined as
$S_{\alpha}(\widehat{d}_{n+\xi|n,x}^{\text{lb}},\widehat{d}_{n+\xi|n,x}^{\text{ub}},d_{n+\xi,x})=(\widehat{d}_{n+\xi|n,x}^{\text{ub}}-\widehat{d}_{n+\xi|n,x}^{\text{lb}})$ $+\frac{2}{\alpha}(\widehat{d}_{n+\xi|n,x}^{\text{lb}}-d_{n+\xi,x})\mathds{1}(d_{n+\xi,x}<\widehat{d}_{n+\xi|n,x}^{\text{lb}})$
$+\frac{2}{\alpha}(d_{n+\xi,x}-\widehat{d}_{n+\xi|n,x}^{\text{ub}})\mathds{1}(d_{n+\xi,x}>\widehat{d}_{n+\xi|n,x}^{\text{ub}}).$
The interval score rewards a narrow prediction interval if and only if the holdout observation lies within the prediction interval. The optimal interval score is achieved when $d_{n+\xi,x}$ not only lies between the $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$, but also the distance between $\widehat{d}_{n+\xi|n,x}^{\text{lb}}$ and $\widehat{d}_{n+\xi|n,x}^{\text{ub}}$ is minimal for a given age $x$.
For different ages and years in the forecasting period, the mean interval score is defined by
$\overline{S}_{\alpha}(h)=\frac{\sum^{20}_{\xi=h}\sum^{111}_{x=1}S_{\alpha}(\widehat{d}_{n+\xi|n,x}^{\text{lb}},\widehat{d}_{n+\xi|n,x}^{\text{ub}};d_{n+\xi,x})}{111\times(21-h)},$
where $S_{\alpha}(\widehat{d}_{n+\xi|n,x}^{\text{lb}},\widehat{d}_{n+\xi|n,x}^{\text{ub}},d_{n+\xi,x})$ denotes the interval score at the $\xi$th curve in the forecasting period. Since we require at least two curves to compute standard deviation, the forecast horizon is $h=1,2,\dots,19$.
4.6 Comparison of interval forecast accuracy
Averaging over the interval forecasts from the 24 countries, Table LABEL:tab:2 presents horizon-specific forecast errors for comparing the accuracy of the 80% prediction intervals constructed by the two transformations. Differing from the point forecast accuracy, the CDF transformation does not consistently outperform the CLR transformation, but it generally provides superior accuracy with smaller CPD and mean interval score. Further, we compare two ways of selecting the number of components in the principal component analysis for modeling the unconstrained data in the transformed space. Due to the limited space, we do not report the results for the 95% nominal coverage probability, but these can be obtained upon request from the corresponding author.
Table 2: Comparison of the interval forecast accuracy, as measured by the CPDα and mean interval score $\overline{S}_{\alpha}$, between the CLR and CDF transformations at the level of significance $\alpha=0.2$. For each transformation, we further compare two ways of selecting the number of components.
Female Male
EVR $K=6$ EVR $K=6$
Metric $h$ CLR CDF CLR CDF CLR CDF CLR CDF
CPDα 1 0.027 0.023 0.089 0.100 0.152 0.047 0.082 0.102
2 0.134 0.040 0.037 0.015 0.143 0.002 0.040 0.010
3 0.006 0.049 0.062 0.042 0.039 0.053 0.005 0.004
4 0.131 0.089 0.097 0.091 0.135 0.116 0.086 0.012
5 0.103 0.070 0.064 0.040 0.052 0.146 0.147 0.000
6 0.103 0.005 0.048 0.047 0.676 0.022 0.014 0.018
7 0.073 0.103 0.045 0.017 0.160 0.025 0.079 0.088
8 0.027 0.030 0.093 0.091 0.027 0.155 0.011 0.008
9 0.040 0.102 0.062 0.093 0.197 0.090 0.008 0.033
10 0.083 0.022 0.002 0.077 0.404 0.185 0.242 0.259
11 0.162 0.029 0.098 0.019 0.178 0.069 0.032 0.080
12 0.179 0.031 0.038 0.043 0.574 0.144 0.043 0.033
13 0.130 0.106 0.153 0.172 0.018 0.026 0.119 0.098
14 0.226 0.115 0.027 0.144 0.096 0.121 0.082 0.134
15 0.117 0.024 0.219 0.064 0.350 0.354 0.159 0.073
16 0.119 0.070 0.040 0.018 0.283 0.187 0.027 0.056
17 0.200 0.117 0.099 0.155 0.160 0.123 0.110 0.186
18 0.197 0.110 0.004 0.083 0.035 0.101 0.052 0.092
19 0.196 0.138 0.133 0.102 0.241 0.151 0.192 0.196
Mean 0.119 0.067 0.074 0.074 0.206 0.112 0.081 0.078
$\overline{S}_{\alpha}$ 1 347 326 149 138 729 415 144 150
2 432 286 241 197 634 339 353 243
3 480 309 333 211 703 699 345 273
4 1301 787 441 404 722 514 392 374
5 263 259 190 151 599 707 337 213
6 621 343 294 305 1846 530 420 428
7 432 321 335 333 999 608 811 578
8 709 575 485 301 2743 1044 706 443
9 770 366 578 355 2461 947 776 717
10 1015 360 395 440 1254 1048 1207 1097
11 1409 1075 975 781 2029 664 979 662
12 2188 527 521 468 2879 1471 1179 610
13 504 645 439 635 2614 918 784 728
14 1432 1803 1514 1702 355 449 471 450
15 1340 601 837 509 2250 1556 1347 1108
16 1313 869 823 479 1999 1641 1137 924
17 1208 472 433 501 2672 1213 1168 975
18 1088 602 518 462 1476 1021 1100 814
19 2824 1500 1429 1556 1985 1884 1789 1516
Mean 1036 633 575 523 1629 930 813 648
While Table LABEL:tab:2 compares the interval forecast accuracy based on the CPD and mean interval scores, aggregated over 24 countries. In Figure 4, we provide several heatmaps that show the number of times one method produces the smallest errors. For modeling the female data, the difference between the CLR and CDF transformations is marginal. For modeling the male data, the CDF transformation with $K=6$ is advocated with the smallest CPD and mean interval score.
(a) CPDα=0.2 for female data
(b) CPDα=0.2 for male data
(c) $\overline{S}_{\alpha=0.2}$ for female data
(d) $\overline{S}_{\alpha=0.2}$ for male data
Figure 4: Heatmaps showing the frequency of the most accurate transformation, as measured by the CPD and mean interval score, at the nominal coverage probability of 80%.
5 Comparisons of life expectancy at birth and annuity price
An immediate application of life-table death count forecasts for individuals aged 60 and over is in the superannuation industry, whose profitability and solvency crucially rely on accurate mortality forecasts to hedge longevity risks appropriately. When a person retires, an optimal way of guaranteeing one’s financial income in retirement is to purchase an annuity product. Such an annuity is a financial contract offered by insurers guaranteeing a steady stream of payments for either a temporary or the lifetime of the annuitant in exchange for an initial premium fee.
Among many annuity products, we study temporary annuities, which have grown in popularity in a number of countries, including Australia and the USA, because lifetime immediate annuities, where rates are locked in for life, have been demonstrated to deliver poor value for money (see, e.g., Cannon & Tonks, 2008, Chapter 6). The temporary annuities pay a pre-determined and guaranteed income level higher than the income provided by a lifetime annuity for a similar amount of premium. Fixed-term annuities also offer an attractive alternative to lifetime annuities and allow the purchaser to buy a deferred annuity later.
With the CDF transformation, we obtain 50-years-ahead forecasts of life-table death counts for all countries. As a vehicle of demonstration, we display the forecasts for Australia in Figure 5 and observe a continuing trend of further longevity. These forecasts are time-series extrapolations of the past trend, and thus should be used with caution. Any extreme shocks, such as COVID-19, are likely to have a substantial impact on the mortality trend.
Figure 5: Using the CDF transformation, we obtain age-specific life-table death count forecasts from 2022 to 2071 for Australian females and males.
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(d_{x})$ (from 2022 to 2071) to their corresponding period life expectancies ($e_{x}$). In Figure 6, we display the forecasted life expectancy at birth of the female and male populations in the 24 countries. In 2022, Japan had the highest life expectancy for females, but by 2050, the highest life expectancy may occur in Switzerland. For the male population, Switzerland had the highest life expectancy in 2022, but by 2050, the highest life expectancy may occur in Australia. Among all the countries, Bulgaria has the lowest life expectancy at birth.
(a) Forecasts for 2022
(b) Forecasts for 2022
(c) Forecasts for 2050
(d) Forecasts for 2050
Figure 6: Estimated period life expectancy at birth of the female and male populations in the 24 countries in 2022 and 2050.
In Figure 7, we display the forecasted life expectancy at age 60 of the female and male populations in the 24 countries. In 2022, Japan had the highest life expectancy for females, but by 2050, the highest life expectancy may occur in Switzerland. For the male population, Australia had the highest life expectancy in 2022, and may still maintain its lead by 2050. Among all the countries, Bulgaria has the lowest life expectancy at age 60.
(a) Forecasts for 2022
(b) Forecasts for 2022
(c) Forecasts for 2050
(d) Forecasts for 2050
Figure 7: Estimated period life expectancy at age 60 of the female and male populations in the 24 countries in 2022 and 2050.
Via a standard life-table calculation, we convert the forecasts of life-table death counts $(d_{x})$ (from 2022 to 2071) to their corresponding survival probabilities ($p_{x}$). We plug the age-specific survival probabilities into the calculation of single-premium term immediate annuities, and we adopt a cohort approach to calculating the survival probabilities, as is the practice of insurance companies, that is, survival probability $p_{60}$ is computed from the forecasted life-table death counts in 2022, $p_{61}$ is computed from the forecasted life-table death counts in 2023, $\dots$, $p_{109}$ is computed from the forecasted life-table death counts in 2071. The $\tau$ year survival probability of a person with entry age $x$ at the contract time is determined by
${}_{\tau}p_{x}=\prod^{\tau}_{j=1}p_{x+j-1}=\prod^{\tau}_{j=1}(1-q_{x+j-1})=\prod^{\tau}_{j=1}(1-\frac{d_{x+j-1}}{l_{x+j-1}}),$
where $d_{x+j-1}$ denotes the number of death counts between ages $x+j-1$ and $x+j$, and $l_{x+j-1}$ represents the number of lives alive at age $x+j-1$.
Once we compute the survival probabilities for ages at and beyond 60, the calculation of annuity prices also depends on the value of the zero-coupon bond price. For an $x$-year-old with a benefit, one monetary amount per year is given by
$a_{x}^{T}=\sum^{T}_{\tau=1}B(0,\tau)_{\tau}p_{x},$
where $B(0,\tau)$ is the $\tau$-year bond price and ${}_{\tau}p_{x}$ denotes the survival probability.
Table 3 provides an example of the annuity calculations using the Australian data. We compute the best estimate of the annuity prices for different entry ages and maturities for female and male policyholders residing in Australia. We assume the constant interest rate at the value of $\lambda=4.1\%$ as of 18 May 2025 and hence the zero-coupon bond is given as $B(0,\tau)=\exp^{-\lambda\tau}$.
Table 3: Australian annuity price estimates for various maturities and initial contract ages. The estimates are based on the 50-years-ahead mortality forecasts and the current interest rate of $4.1\%$ as of 18 May 2025.
Female Male
Age $T=5$ 10 15 20 25 30 5 10 15 20 25 30
60 4.379 7.859 10.589 12.688 14.234 15.289 4.347 7.746 10.355 12.303 13.681 14.573
65 4.359 7.780 10.409 12.345 13.667 14.443 4.314 7.627 10.100 11.849 12.980 13.579
70 4.322 7.644 10.091 11.761 12.741 13.187 4.264 7.447 9.698 11.154 11.925 12.190
75 4.263 7.403 9.547 10.805 11.377 11.527 4.187 7.147 9.062 10.076 10.425 10.479
80 4.135 6.958 8.615 9.369 9.566 9.586 4.031 6.6398 8.020 8.496 8.569 8.572
85 3.939 6.252 7.304 7.579 7.607 3.807 5.822 6.516 6.622 6.628
90 3.552 5.167 5.589 5.632 3.307 4.445 4.6195 4.629
95 3.020 3.809 3.890 2.504 2.888 2.9076
100 2.113 2.328 1.581 1.664
105 1.312 1.003
Figure 8 presents dotcharts for the estimated annuity prices of the 24 countries at entry age 60 with a maturity period of 5 years. With a return of one monetary amount, the estimated annuity price is the highest for Japan, while the two with the lowest values are Iceland and Hungary.
Figure 8: With a return of one monetary amount annually, estimated annuity prices of the female and male populations in the 24 countries at entry age 60 with a maturity period of 5 years.
In Figure 9, we present dotcharts for the estimated annuity prices of the 24 countries at entry age 70 with a maturity period of 20 years. With a return of one monetary amount, the estimated annuity price is relatively the highest for Japan, while the lowest is for Hungary.
Figure 9: With a return of one monetary amount annually, estimated annuity prices of the female and male populations in the 24 countries at entry age 70 with a maturity period of 20 years.
To measure forecast uncertainty, we construct the 50-years-ahead pointwise prediction intervals for the life-table death counts. For each of the 50 years, we search for the optimal tuning parameter that achieves the smallest CPD between the empirical coverage probability and nominal coverage probabilities of 80% and 95%. In Table LABEL:tab:4, we present the 80% and 95% pointwise prediction intervals of annuities for different entry ages and maturities, where age + maturity $\leq 110$.
Table 4: With a return of \$1 annually, the 80% and 95% pointwise prediction intervals of annuity prices with different entry ages and maturities $(T)$ for female and male populations residing in Australia.
Sex Age $T=5$ 10 15 20 25 30
80% nominal coverage probability
F 60 (4.373, 4.386) (7.834, 7.887) (10.530, 10.659) (12.577, 12.822) (14.055, 14.457) (15.044, 15.603)
65 (4.348, 4.372) (7.734, 7.835) (10.306, 10.536) (12.162, 12.578) (13.405, 14.009) (14.118, 14.874)
70 (4.301, 4.349) (7.566, 7.742) (9.923, 10.307) (11.501, 12.105) (12.407, 13.191) (12.818, 13.687)
75 (4.232, 4.304) (7.286, 7.557) (9.330, 9.837) (10.504, 11.215) (11.037, 11.844) (11.185, 11.994)
80 (4.083, 4.203) (6.816, 7.149) (8.386, 8.929) (9.098, 9.742) (9.296, 9.936) (9.321, 9.946)
85 (3.898, 3.996) (6.137, 6.411) (7.153, 7.513) (7.435, 7.776) (7.472, 7.791)
90 (3.506, 3.615) (5.097, 5.264) (5.537, 5.657) (5.595, 5.679)
95 (2.990, 3.041) (3.703, 3.883) (3.743, 3.993)
100 (1.975, 2.203) (2.086, 2.490)
105 (0.858, 1.543)
M 60 (4.336, 4.359) (7.699, 7.799) (10.253, 10.477) (12.139, 12.505) (13.452, 13.972) (14.275, 14.965)
65 (4.292, 4.341) (7.550, 7.722) (9.956, 10.282) (11.632, 12.133) (12.682, 13.387) (13.212, 14.097)
70 (4.235, 4.300) (7.363, 7.557) (9.541, 9.912) (10.907, 11.507) (11.596, 12.410) (11.8218, 12.741)
75 (4.165, 4.216) (7.065, 7.265) (8.884, 9.330) (9.801, 10.499) (10.102, 10.927) (10.147, 10.994)
80 (3.987, 4.096) (6.488, 6.869) (7.748, 8.441) (8.162, 9.015) (8.225, 9.1051) (8.228, 9.110)
85 (3.730, 3.922) (5.610, 6.144) (6.228, 6.956) (6.321, 7.083) (6.326, 7.090)
90 (3.209, 3.447) (4.263, 4.707) (4.422, 4.904) (4.430, 4.914)
95 (2.457, 2.565) (2.828, 2.966) (2.847, 2.987)
100 (1.569, 1.596) (1.650, 1.680)
105 (1.000, 1.006)
95% nominal coverage probability
F 60 (4.369, 4.392) (7.818, 7.911) (10.494, 10.717) (12.513, 12.931) (13.957, 14.635) (14.916, 15.849)
65 (4.341, 4.383) (7.709, 7.878) (10.250, 10.636) (12.068, 12.758) (13.274, 14.270) (13.962, 15.199)
70 (4.289, 4.370) (7.525, 7.818) (9.840, 10.472) (11.377, 12.363) (12.253, 13.524) (12.650, 14.055)
75 (4.216, 4.335) (7.230, 7.671) (9.232, 10.047) (10.372, 11.507) (10.890, 12.175) (11.036, 12.326)
80 (4.061, 4.250) (6.757, 7.276) (8.293, 9.136) (8.990, 9.986) (9.186, 10.178) (9.214, 10.185)
85 (3.882, 4.029) (6.095, 6.504) (7.098, 7.636) (7.381, 7.892) (7.421, 7.902)
90 (3.490, 3.650) (5.072, 5.318) (5.518, 5.696) (5.581, 5.710)
95 (2.975, 3.048) (3.649, 3.908) (3.674, 4.028)
100 (1.902, 2.233) (1.971, 2.544)
105 (0.622, 1.610)
M 60 (4.328, 4.369) (7.671, 7.842) (10.193, 10.574) (12.047, 12.666) (13.330, 14.203) (14.124, 15.279)
65 (4.279, 4.362) (7.509, 7.794) (9.882, 10.421) (11.524, 12.352) (12.541, 13.703) (13.043, 14.506)
70 (4.221, 4.327) (7.323, 7.640) (9.470, 10.074) (10.798, 11.779) (11.454, 12.791) (11.666, 13.175)
75 (4.156, 4.238) (7.031, 7.353) (8.811, 9.533) (9.689, 10.829) (9.973, 11.320) (10.015, 11.397)
80 (3.970, 4.143) (6.428, 7.044) (7.641, 8.767) (8.033, 9.420) (8.091, 9.523) (8.094, 9.528)
85 (3.699, 4.010) (5.525, 6.391) (6.114, 7.294) (6.203, 7.436) (6.207, 7.443)
90 (3.169, 3.544) (4.191, 4.886) (4.345, 5.098) (4.353, 5.109)
95 (2.439, 2.599) (2.806, 3.010) (2.824, 3.031)
100 (1.565, 1.602) (1.645, 1.687)
105 (0.999, 1.007)
The forecast uncertainties increase in prediction interval width as maturities increase from $T=5$ to $T=30$ for a given age. They also increase as the initial ages when entering contracts increase from 60 to 105 for a given maturity.
Figures 10a and 10b show the 80% and 95% prediction intervals for annuity prices at entry age 60 with 5-year maturity across 24 countries, while Figures 10c and 10d show the results for entry age 70 with 20-year maturity. With a return of one monetary unit, estimates are shown as open circles, with open intervals indicating prediction intervals. Forecast uncertainty is noticeably higher for Iceland and lower for Hungary. With increasing maturity, prediction intervals widen, reflecting greater long-term uncertainty.
(a) Entry age 60 with 5-year maturity
(b) Entry age 60 with 5-year maturity
(c) Entry age 70 with 20-year maturity
(d) Entry age 70 with 20-year maturity
Figure 10: With a return of one monetary amount annually, the 80% and 95% pointwise prediction intervals of the annuity prices of the female and male populations in the 24 countries.
6 Conclusion
For modeling and forecasting age distribution of death counts, we consider the centered log-ratio and cumulative distribution function transformations. Both transformations allow us to bring a constrained object to an unconstrained space, where the linear techniques can be applied. Such a technique is principal component analysis, which can summarize and model variation in the unconstrained data. Using 24 countries’ data, we evaluate the point forecast accuracy (measured by the Kullback-Leibler and Jensen-Shannon divergences) and the interval forecast accuracy (measured by the coverage probability difference between empirical and nominal coverage probabilities and mean interval score). For producing both point and interval forecasts, the cumulative distribution function transformation provides smaller point and interval forecast errors than the centered log-ratio transformation. Between the two ways of selecting the number of principal components, we advocate the selection of $K=6$ for forecasting.
We apply the cumulative distribution function transformation to generate 50-years-ahead forecasts of age-specific life-table death counts for 24 countries. From the forecasted age distribution of death counts, we evaluate and compare the life expectancy at birth across multiple countries. From the forecasted life-table death counts, we convert them into age-specific survival probabilities and obtain temporary annuity prices. As expected, we observe that the survival probability has a pronounced impact on annuity prices. Coupling with the uncertainty in interest rates, we quantify the uncertainty of annuity prices via pointwise prediction intervals. For reproducibility, the computer [Uncaptioned image] code is available at https://github.com/hanshang/CLR_vs_CDF_transformation.
There are a few ways in which the methodology presented in the paper could be extended, and we briefly discuss two. First, although we calculate temporary annuities, other types of annuity prices, such as the whole-life immediate annuity or deferred annuity, are possible. Second, we could consider cohort life-table death counts for modeling a group of individuals, such as baby boomers.
Acknolwedgment
This research was supported by the Australian Research Council Future Fellowship (grant no. FT240100338).
Disclosure statement
No potential conflict of interest was reported by the authors.